the number of red poinsettia plants is 25 more than 7 time the number of white ones. there are 225 total. how many are white.
Answer by macston(5194)

.
W=number of white; R=number of red=7W+25
.
R+W=225
(7W+25)+W=225
8W=200
W=25
ANSWER: There are 25 white poinsettias.
.
CHECK:
R=7W+25
R=7(25)+25
R=175+25
R=200
There are 200 red poinsettias.
.
R+W=225
200+25=225
225=225